OpenStudy (anonymous):

Write - 23/4 as a mixed number

jhonyy9 (jhonyy9):

-64/8

jhonyy9 (jhonyy9):

or (-2)*(23/8)

jhonyy9 (jhonyy9):

-44/8 - 1/4 = (-44-2)/8 = -46/8 = -23/4